# Inference on many jumps in nonparametric panel regression models

### v2

@@ 3.2 Gaussian Approximation Results
3.3 Asymptotic Validity of the Feasible Testing Procedures
@@
Theorem 3.3 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.6 hold and $N(\underline{b}T)^{-q/4}\rightarrow 0.$ Then we have $\max_{j\in[N]}|\hat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N))$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N)).$
@@
Theorem 3.4 (Power).
Suppose that Assumptions 3.1-3.6 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
@@
$(ii)$ If $\mathcal{R}_{NT}\to 0$ in Theorem 3.2 and $\max_{j\in\left[N\right]}(Tb_{j})^{1/2}|\gamma_{j}-\bar{\gamma}|\gg(\mathrm{log}N)^{1/2},$ then $\lim_{\left(N,T\right)\rightarrow\infty}\mathbb{P}(\mathcal{\hat{Q}}>q_{\alpha})=1.$
@@
3.4 Theoretical Results for the Unknown Threshold Case
In this section, we present results related to unknown thresholds. The following theorem provides theoretical support for the GA in cases with unknown $c_{0j}$’s. For simplicity of notation, we define $\mathcal{C}_{j}$ as the unified interval $\mathcal{C}_{j}=[c_{\min},c_{\max}]$, where $c_{\min}$ and $c_{\max}$ are constants in $\mathbb{R}$. The theorem still holds even when $\mathcal{C}_{j}$ differs for different $j$. Recall that $[c_{1},\ldots,c_{K}]$ is the grid we search over. Let $c_{1}=c_{\min}$ and $c_{K}=c_{\max}.$ As our statistics aggregate over different threshold grids, the problem becomes more demanding. Let $\Delta_{c,\min}=\min_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$ and $\Delta_{c,\max}=\max_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$.
and $\underline{d}^{C}=(d_{1K}^{\top},d_{2K}^{\top},\ldots,d_{NK}^{\top})^{\top}.$ Define $Z^{C}=\left({{Z_{1}^{C\top},...,Z_{N}^{C\top}}}\right)^{\top}$ as an $NK$-vector of mean zero normally distributed random variables with variance-covariance $\Sigma^{C}$ specified below. For each $j\in[N]$ and $1\leq i_{1},i_{2}\leq K$, the covariance between the $i_{1}$- and $i_{2}$-th elements of $Z_{j}^{C}$ is given by
Theorem 3.5.
(Gaussian Approximation) Suppose that Assumptions 3.1-3.6 hold. Suppose that $\Delta_{c,\max}\rightarrow 0.$ Then
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{(NK)T}.$ (3.8)
The proof of the above theorem is a generalization of Theorem 3.1 with a different variance-covariance structure. It should be noted that the conclusion of the above theorem follows the same pattern as Theorem 3.1, except that $N$ is replaced with $NK$. The following theorem provides the theoretical support for the consistency of the threshold estimator. Note that we only focus on the location $j$ where $\gamma_{j}$ is not equal to 00, and we require those breaks to be significant.
Theorem 3.6 (Consistency of the estimation of thresholds).
Suppose that Assumptions 3.1-3.6 hold. Assume $\min_{1\leq j\leq N:\gamma_{j}\neq 0}(Tb_{j})^{1/2}|\gamma_{j}|\gg 1$ and $\Delta_{c,\max}\rightarrow 0.$ Then
Under the assumption of a minimum break signal, we can achieve consistency in detecting the break. The rate is expected to depend on both the signal strength $\gamma_{j}$ and the available sample size $T$. Note that the precision of threshold estimation is not affected by the bandwidth $b_{j}$. This rate is in line with the high-dimensional changepoint literature, see for example Theorem 2 in Li et al., (2024).

### v3

@@ A.2 A Homogeneity Test
A.3 Asymptotic Validity of the Feasible Testing Procedures
@@
Theorem A.2 (Consistency of the variance estimator).
Suppose that Assumptions 3.1- 3.3 and SA-1-SA-3 hold and $N(\underline{b}T)^{-q/4}\rightarrow 0.$ Then we have $\max_{j\in[N]}|\hat{\sigma}_{e,j}^{2}-\sigma_{e,j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N))$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}((\underline{b}T)^{-1/2}\mathrm{log}(N)).$
@@
Theorem A.3 (Power).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$
@@
$(ii)$ If $\mathcal{R}_{NT}\to 0$ in Theorem A.1 and $\max_{j\in\left[N\right]}(Tb_{j})^{1/2}|\gamma_{j}-\bar{\gamma}|\gg(\mathrm{log}N)^{1/2},$ then $\lim_{\left(N,T\right)\rightarrow\infty}\mathbb{P}(\mathcal{\hat{Q}}>q_{\alpha})=1.$
@@
A.4 Theorems for the Unknown Threshold Case
$\hat{\sigma}_{e,j}^{2}(c_{i})=\frac{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}\phi_{\acute{a}}(\hat{e}_{jt}^{2})}{\sum_{t=1}^{T}\mathbf{1}_{\{|X_{jt}-c_{i}|\leq b_{j}\}}}.$ (SA-4)
Theorem A.4 (Consistency of the variance estimator).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold and $N(bT)^{-q/4}\rightarrow 0.$ Then we have
$\max_{j\in[N]}\max_{1\leq i\leq K}|\hat{\sigma}_{e,j}^{2}(c_{i})-\sigma_{e,j}^{2}(c_{i})|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$ and $\max_{j\in[N]}|\hat{v}_{j}^{2}-v_{j}^{2}|$ $=O_{\mathbb{P}}(\mathrm{log}^{1/2}(NK)\acute{a}b^{*}/\bar{b}+\acute{a}^{-p/2+1}+b^{*2})$.
Theorem A.5 (Consistency of the tests).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Let $\bar{\gamma}=\sum_{j=1}^{N}\gamma_{j}/N.$